Explain the working of refridgerator  and obtain an expression for its coffecient of performance.

Grade:12

1 Answers

Ashutosh Sharma
7 years ago
Hi.Refrigerator Is basically Works on the principle of Heat Pump. It can be defined as reverse function of heat engine. The heat is transmitted sink to source. Whereas, The Temp Of Sink is higher And Source is having lower temperature. Thus, in the case of heat Pump or refrigerator, we have to do some work. Providing electricity is work here. The coefficent of performance is the ratio of Desired Result to the work. W=Q1-Q2COP=Q2/W COP=Q2/Q2-Q1